# Nonstrange and strange quark matter at finite temperature within modified NJL model and protoquark stars

### v1

@@ III Quark matter in $\beta$-equilibrium
In this section, we investigate the effects of exchange interactions and temperature on the properties of quark matter under conditions of $\beta$-equilibrium and charge neutrality, neglecting neutrino contributions. Figure 1 presents the effective quark mas and quark number densities of u,d,su,\ d,\ s quarks, as well as electron chemical potential calculated within the $(2+1)$-flavor and two-flavor NJL models. At zero temperature, when $\tilde{\mu}_{f}<M_{f}$, the quark condensate is independent of the quark chemical potential, and the dynamical quark mass remains at its vacuum value. When $\tilde{\mu}_{f}>M_{f}$, the chiral symmetry begins to restore. As the parameter $\alpha$ becomes nonzero, vector interactions are introduced self-consistently into the system. Due to the repulsive nature of these interactions in the exchange channel, the dynamical quark mass decreases more gradually with increasing chemical potential $\mu$, particularly for the $s$ quark. This effect is clearly illustrated by comparing the cases of $\alpha=0.0,\ T=0.00$ GeV and $\alpha=0.5,\ T=0.00$ GeV. A further increase of $\alpha$ to 0.8 (0.9) enhances the role of vector interactions in the exchange channels, further suppresses chiral symmetry restoration, and drives the first-order chiral phase transition toward a crossover. In comparison with the NJL model without vector interactions ($\alpha=0.0,\ T=0.00$ GeV), increasing the temperature makes the quark condensate begins to restore at a relatively lower chemical potential.
@@
As expected, at zero temperature, the quark number density stays zero when $\mu$ is smaller than the constituent quark mass $M_{u,d}$ and $M_{s}$. Once $\mu$ exceeds the corresponding dynamical quark mass, the quark number density exhibits a discontinuous jump to a nonzero value, signaling a first-order phase transition. In the two-flavor quark matter system, the restoration of the quark condensate leads to an increase in the up and down quark number densities, $\rho_{u,d}$. Due to the charge neutrality condition, this necessitates an increase in the electron fraction, and thus, the electron chemical potential $\mu_{e}$ rises with increasing $\mu$. In contrast, the behavior of $\mu_{e}$ in strange quark matter is more intricate. Increase the value of $\alpha$ enhances the vector interaction strength, and these repulsive interactions suppress the restoration of the strange quark condensate, thereby reducing the strange quark number density $\rho_{s}$. Consequently, particularly for $\mu\gtrsim 400$ MeV, a larger value of $\alpha$ leads to a higher electron chemical potential, compensating for the reduced negative charge from strange quarks in order to maintain overall charge neutrality. In particular, for $\alpha=0.9$, the strange quark condensate is rarely restored, yielding a system with a negligible strange quark population and, consequently, a significantly elevated $\mu_{e}$ at high $\mu$.
@@
To investigate the properties of proto-quark stars, we assume that neutrinos are trapped in the matter. In the following analysis, we consider a temperature range of $T=0-40\,\mathrm{MeV}$ and a total lepton fraction of $Y_{l}=0.1-0.4$, consistent with conditions expected in compact stars during the core-collapse phase of massive stellar evolution Burrows & Lattimer (1986). We will demonstrate that the strength of exchange interactions and the lepton fraction plays a more significant role than temperature in determining the stiffness of the EOS. The underlying reasons for this behavior will be discussed later.
@@
For neutrino-trapped matter, we solve Eqs. (4 or 14), (22), (24), and (25) self-consistently, taking the chemical potential of the $u$ quark as the input quantity. The resulting physical solutions for two-flavor and (2+1)-flavor quark matter, corresponding to a representative case with $\alpha=0.5$, $Y_{l}=0.1$, and $T=0.04~\mathrm{GeV}$, are shown in Fig. 2. For (2+1)-flavor quark matter, When $\mu>220~\mathrm{MeV}$, the $u$ quark condensate begins to melt slightly, accompanied by the appearance of a small population of $u$ and $d$ quarks. Owing to flavor-mixing effects and the associated reduction of $M_{u}$ and $M_{d}$, the dynamical mass of the $s$ quark also decreases slightly. At $\mu=351~\mathrm{MeV}$, a first-order phase transition occurs, marked by a discontinuous drop in the dynamical masses $M_{u}$, $M_{d}$, and $M_{s}$. The effective $u$ quark mass differs slightly from the $d$ quark mass due to the $\beta$-equilibrium. For $\mu>365~\mathrm{MeV}$, the $ud$ quark masses are fully restored to their current masses, and the flavor-mixing effect becomes negligible. As the quark chemical potential increases further, the strange quark condensate continues to decrease, and the number density of $s$ quarks increases. In contrast, the two-flavor case exhibits a relatively low dynamical quark mass of approximately $311~\mathrm{MeV}$, generated via spontaneous chiral symmetry breaking, with only a weak discontinuity in the $M(\mu)$ relation. Due to the absence of strange quarks, both the electron chemical potential $\mu_{e}$ and the neutrino chemical potential $\mu_{\nu_{e}}$ increase monotonically with $\mu_{u}$, driven by the requirements of charge neutrality and $\beta$-equilibrium.
@@
In Fig. 3, we examine the effects of lepton fraction $Y_{l}$, temperature $T$, and the strength of exchange interactions—characterized by the parameter $\alpha$—on both strange and nonstrange quark matter. We present the effective quark masses $M_{u}$ and $M_{s}$ as functions of the quark chemical potential $\mu$, along with the pressure as a function of baryon number density $\rho_{B}$ for various parameter sets. In the lower panel of Fig. 3, for the first-order chiral phase transition, we consider only the chirally restored quark matter phase predicted by the NJL model to be reliable; therefore, the results are shown starting from $2\rho_{0}$. Under the same conditions of $Y_{l}=0.1$ and $T=0.04$ GeV, comparing stronger vector interactions with $\alpha=0.5$ to those with $\alpha=0.0$ shows that the former shifts the onset of the first-order phase transition to higher quark chemical potentials and significantly suppresses the restoration of effective quark masses at high chemical potentials, especially that of the strange quark. A similar trend is observed in the case of nonstrange quark matter. As shown in the lower panel, the EOS becomes stiffer for larger $\alpha$, reflecting a stronger contribution from vector interactions in the exchange channels, for both strange and nonstrange quark matter. Compared to the case of $\alpha=0.8$, $Y_{l}=0.1$, and $T=0.01$ GeV, increasing the temperature to $T=0.04$ GeV accelerates the melting of the dynamical quark masses before $\mu\sim 400\;\text{MeV}$, causing the strange quark phase to emerge at lower chemical potentials. The similar behavior can also be found for nonstrange quark matter. The temperature has only a slight effect on the stiffness of the EOS, as can be seen from the $P(\rho_{B})$ relations. For fixed $\alpha=0.8$ and $T=0.04$ GeV, increasing the lepton fraction from $Y_{l}=0.1$ to $Y_{l}=0.4$ suppresses the restoration of dynamical masses and stiffens the EOS of quark matter. A higher lepton fraction implies a larger electron population, which, due to the charge neutrality condition, suppresses the abundance of other negatively charged particles, such as the $s$ quark. As a result, a higher $Y_{l}$ leads to a reduced fraction of $s$ quarks, thereby hindering the restoration of their dynamical mass.
Figure 3: The effective quark masses of u,d,su,\ d,\ s quarks versus quark chemical potential $\mu$ for various choices of the (2+1)-flavor and two-flavor NJL model parameters with neutrino trapped scenario for the study of their effects (see text for details). Lower panel: The pressure as a function of baryon number density, expressed in units of the nuclear saturation density $\rho_{0}$, for both strange quark matter and nonstrange quark matter.
V Proto-quark star properties
[…]
Figure 5: Left panel: The pressure versus the energy density (EOS) for several parameter sets calculated from the two-flavor modified NJL model. The EOSs are all smooth here. Right panel: Corresponding mass-radius relations for the self-bound nonstrange quark matter EOS for the six parameter sets shown in the left panel. The available astronomical constraints shown here are the same as in Fig. 4.
In the following, we examine the influence of the exchange interaction parameter $\alpha$, lepton fraction $Y_{l}$, and temperature $T$ on the EOS and the structure of protoquark stars. As shown in Sec. IV, increasing the strength of the exchange channel $\alpha$ alters the nature of the chiral phase transition from first-order to a crossover. Since the relationship between chiral symmetry restoration and deconfinement remains unclear, in this work, we assume that deconfinement occurs simultaneously with the chiral phase transition. Therefore, to effectively mimic confinement at low densities, a finite vacuum pressure is introduced. According to the properties of the chiral phase transition, we adopt different prescriptions for defining the vacuum pressure B-B:
  • For a first-order phase transition at low temperature and weak vector interaction, the chiral symmetry breaking phase (hadronic phase), which cannot be reliably described by the NJL model, will be transformed to the chiral restored phase (quark phase). Therefore, we consider only the quark phase after the transition and choose the minimum vacuum pressure $B_{0}$ such that the pressure vanishes at a finite energy density, as required for modeling self-bound quark stars. Any other choice of $B$ should be larger than $B_{0}$. For example, in Fig. 4, for the case of $\alpha=0.5$, $Y_{l}=0.4$, and $T=40~\mathrm{MeV}$, the chiral phase transition is of first order, resulting in a discontinuity in the energy density of the EOS. To ensure that the quark star is composed of chirally restored quark matter, the vacuum pressure must exceed $133,\mathrm{MeV}$, allowing the quark matter EOS to reach zero pressure at a finite energy density—thus ensuring the star is self-bound at the surface.
  • For a crossover chiral transition at relatively large temperature and strong vector interactions in the exchange channel, the transition is smooth. As a result, the choice of vacuum pressure is more flexible, and even a small vacuum pressure can ensure that the EOS yields a finite energy density at zero pressure.
@@ V Proto-quark star properties
We note that, according to the Bodmer–Witten hypothesis Bodmer (1971); Witten (1984), absolutely stable quark matter must have an energy per baryon lower than that of the $^{56}\mathrm{Fe}$ nucleus at zero pressure and zero temperature. Accordingly, for self-bound strange quark matter, the condition $(E/A){\mathrm{uds}}\leq 930~\mathrm{MeV}$ must be satisfied for the hypothesis of strange matter to hold. In contrast, bulk nonstrange quark matter should have an energy per baryon higher than that of the confined phase, i.e., $(E/A){\mathrm{ud}}\geq 930~\mathrm{MeV}$. For absolutely stable nonstrange quark matter, the stability condition is reversed: $(E/A){\mathrm{ud}}\leq 930~\mathrm{MeV}\leq(E/A){\mathrm{uds}}$. In our previous work using the modified NJL model with exchange interactions Yuan et al. (2022), we demonstrated that there is sufficient parameter space to support both absolutely stable nonstrange and strange quark matter at zero temperature. However, at finite temperature, these conditions are not necessarily required to be satisfied.
V.1 Strange proto-quark star properties
@@
In Fig. 4, we present the EOS at finite temperature with neutrino trapping, along with the corresponding mass-radius relations for protoquark stars. The full EOSs represent results derived from the NJL model, including the chirally broken phase. To model self-bound quark matter that can form quark stars, a vacuum pressure $B$ is added to ensure the system reaches zero pressure at a finite density. For the case of $\alpha=0.5$, $Y_{l}=0.1$, and $T=10\,\mathrm{MeV}$, a discontinuous jump in energy density appears, characteristic of a first-order phase transition. In this scenario, the vacuum pressure must be at least $B_{0}=133\,\mathrm{MeV}$ to ensure that the pressure vanishes at the surface of the compact star. However, when the chiral phase transition becomes a crossover, the required vacuum pressure $B$ can be smaller than $B_{0}$.

### v2

@@ III Quark matter in $\beta$-equilibrium
In this section, we investigate the effects of exchange interactions and temperature on the properties of quark matter under conditions of $\beta$-equilibrium and charge neutrality, neglecting neutrino contributions. Figure 1 presents the constituent quark masses and quark number densities of u,d,su,\ d,\ s quarks, as well as electron chemical potential calculated within the three-flavor and two-flavor NJL models. At zero temperature, when $\tilde{\mu}_{f}<M_{f}$, the quark condensate is independent of the quark chemical potential, and the constituent quark mass remains at its vacuum value. When $\tilde{\mu}_{f}>M_{f}$, the chiral symmetry begins to restore. As the parameter $\alpha$ becomes nonzero, vector interactions are introduced self-consistently into the system. Due to the repulsive nature of these interactions in the exchange channel, the constituent quark mass decreases more gradually with increasing chemical potential $\mu$, particularly for the $s$ quark. This effect is clearly illustrated by comparing the cases of $\alpha=0.0,\ T=0.00$ MeV and $\alpha=0.5,\ T=0.00$ MeV. A further increase of $\alpha$ to 0.8 (0.9) enhances the role of vector interactions in the exchange channels, further suppresses chiral symmetry restoration, and drives the first-order chiral phase transition toward a crossover. In comparison with the NJL model without vector interactions ($\alpha=0.0,\ T=0.00$ MeV), increasing the temperature makes the quark condensate begins to restore at a relatively lower chemical potential.
@@
As expected, at zero temperature, the quark number density stays zero when $\mu$ is smaller than the constituent quark mass $M_{u,d}$ and $M_{s}$. Once $\mu$ exceeds the corresponding constituent quark mass, the quark number density exhibits a discontinuous jump to a nonzero value, signaling a first-order phase transition. In the two-flavor quark matter system, the restoration of the quark condensate leads to an increase in the up and down quark number densities $\rho_{u,d}$. Due to the charge neutrality condition, this necessitates an increase in the electron fraction, and thus, the electron chemical potential $\mu_{e}$ rises with increasing $\mu$. In contrast, the behavior of $\mu_{e}$ in strange quark matter is more intricate. Increase the value of $\alpha$ enhances the vector interaction strength, and these repulsive interactions suppress the restoration of the strange quark condensate, thereby reducing the strange quark number density $\rho_{s}$. Consequently, particularly for $\mu\gtrsim 400$ MeV, a larger value of $\alpha$ leads to a higher electron chemical potential, compensating for the reduced negative charge from strange quarks in order to maintain overall charge neutrality. In particular, for $\alpha=0.8$, the strange quark condensate is rarely restored, yielding a system with a negligible strange quark population and, consequently, a significantly elevated $\mu_{e}$ at high $\mu$.
Figure 3: The constituent quark masses of u,d,su,\ d,\ s quarks versus quark chemical potential $\mu$ for various choices of the three- and two-flavor NJL model parameters with neutrino trapped scenario for the study of their effects (see text for details). Lower panel: The pressure as a function of baryon number density, expressed in units of the nuclear saturation density $\rho_{0}$, for both strange quark matter and nonstrange quark matter.
@@
To investigate the properties of protoquark stars, we assume that neutrinos are trapped in the matter. In the following analysis, we consider a temperature range of $T=0-40\;\text{MeV}$ and a total lepton fraction of $Y_{l}=0.1-0.4$, consistent with conditions expected in compact stars during the core-collapse phase of massive stellar evolution Burrows & Lattimer (1986). We will demonstrate that the strength of exchange interactions and the vacuum pressure play a more significant role than the lepton fraction and temperature in determining the stiffness of the EOS. The underlying reasons for this behavior will be discussed later.
@@
For neutrino-trapped matter, we solve Eqs. (4 or 13), (21), (23), and (24) self-consistently, taking the chemical potential of the $u$ quark as the input quantity. The resulting physical solutions for two- and three-flavor quark matter, corresponding to a representative case with $\alpha=0.5$, $Y_{l}=0.1$, and $T=40\;\text{MeV}$, are shown in Fig. 2. For three-flavor quark matter, When $\mu>220~\mathrm{MeV}$, the $u$ quark condensate begins to melt slightly, accompanied by the appearance of a small population of $u$ and $d$ quarks. Owing to flavor-mixing effects and the associated reduction of $M_{u}$ and $M_{d}$, the constituent mass of the $s$ quark also decreases slightly. At $\mu=351~\mathrm{MeV}$, a first-order phase transition occurs, marked by a discontinuous drop in $M_{u}$, $M_{d}$, and $M_{s}$. The $M_{u}$ differs slightly from the $d$ quark mass due to the $\beta$-equilibrium. For $\mu>365~\mathrm{MeV}$, the $ud$ quark masses are fully restored to their current masses, and the flavor-mixing effect becomes negligible. As the quark chemical potential increases further, the strange quark condensate continues to decrease, and the number density of $s$ quarks increases. In contrast, the two-flavor case exhibits a relatively low constituent quark mass of approximately $311~\mathrm{MeV}$, generated via spontaneous chiral symmetry breaking, with only a weak discontinuity in the $M(\mu)$ relation. Due to the absence of strange quarks, both the electron chemical potential $\mu_{e}$ and the neutrino chemical potential $\mu_{\nu_{e}}$ increase monotonically with $\mu_{u}$, driven by the requirements of charge neutrality and $\beta$-equilibrium.
@@
In Fig. 3, we examine the effects of lepton fraction $Y_{l}$, temperature $T$, and the strength of exchange interactions–characterized by the parameter $\alpha$–on both strange and nonstrange quark matter. We present the constituent quark masses $M_{u}$ and $M_{s}$ as functions of the quark chemical potential $\mu$, along with the pressure as a function of baryon number density $\rho_{B}$ for various parameter sets. In the lower panel of Fig. 3, for the first-order chiral phase transition, we consider only the chirally restored quark matter phase predicted by the NJL model to be reliable; therefore, the results are shown starting from $\rho_{0}$. Under the same conditions of $Y_{l}=0.1$ and $T=40\;\text{MeV}$, comparing stronger vector interactions with $\alpha=0.5$ to those with $\alpha=0.0$ shows that the former shifts the onset of the first-order phase transition to higher quark chemical potentials and significantly suppresses the restoration of constituent quark masses at high chemical potentials, especially that of the strange quark. A similar trend is observed in the case of nonstrange quark matter. As shown in the lower panel, the EOS becomes stiffer for larger $\alpha$, reflecting a stronger contribution from vector interactions in the exchange channels. Compared to the case of $\alpha=0.8$, $Y_{l}=0.1$, and $T=10\;\text{MeV}$, increasing the temperature to $T=40\;\text{MeV}$ accelerates the melting of the constituent quark masses before $\mu\sim 400\;\text{MeV}$, causing the strange quark phase to emerge at lower chemical potentials. This behavior can also be found for nonstrange quark matter. The temperature has only a slight effect on the stiffness of the EOS, as can be seen from the $P(\rho_{B})$ relations. For fixed $\alpha=0.8$ and $T=40\;\text{MeV}$, increasing the lepton fraction from $Y_{l}=0.1$ to $Y_{l}=0.4$ slightly suppresses the restoration of constituent masses and induces a modest change in the stiffness the EOS of quark matter. A higher lepton fraction implies a larger electron population, which, due to the charge neutrality condition, suppresses the abundance of other negatively charged particles, such as the $s$ quark. As a result, a higher $Y_{l}$ leads to a reduced fraction of $s$ quarks, thereby hindering the restoration of their constituent mass. A analogous effect was also reported in Ref. Yasutake & Kashiwa (2009), which investigated the effects of leptons on proto-neutron stars.
Figure 4: Upper panel: EOS for several representative parameter sets calculated within NJL model. Results are shown for the three-flavor case with ($\alpha=0.5$, $Y_{l}=0.1$, $T=40$ MeV, $B^{1/4}=133\;\text{MeV}$), ($\alpha=0.8$, $Y_{l}=0.1$, $T=40$ MeV, $B^{1/4}=133\;\text{MeV}$), ($\alpha=0.8$, $Y_{l}=0.1$, $T=10$ MeV, $B^{1/4}=133\;\text{MeV}$), ($\alpha=0.8$, $Y_{l}=0.1$, $T=10$ MeV, $B^{1/4}=95\;\text{MeV}$), ($\alpha=0.8$, $Y_{l}=0.4$, $T=10$ MeV, $B^{1/4}=95\;\text{MeV}$), as well as for the two-flavor case with ($\alpha=0.8$, $Y_{l}=0.1$, $T=10$ MeV, $B^{1/4}=110\;\text{MeV}$), ($\alpha=0.8$, $Y_{l}=0.1$, $T=40$ MeV, $B^{1/4}=110\;\text{MeV}$), ($\alpha=0.8$, $Y_{l}=0.1$, $T=10$ MeV, $B^{1/4}=90\;\text{MeV}$), ($\alpha=0.8$, $Y_{l}=0.4$, $T=10$ MeV, $B^{1/4}=90\;\text{MeV}$), ($\alpha=0.9$, $Y_{l}=0.1$, $T=10$ MeV, $B^{1/4}=90\;\text{MeV}$). The panel shows the full EOS, featuring the phase with unrestored chiral symmetry on the left and the restored quark phase corresponding to self-bound strange quark matter with a nonzero vacuum pressure $B$ on the right. Middle panel: Free energy per baryon $f/\rho_{B}$ as a function of pressure $P$, and the squared speed of sound $c_{s}^{2}$ as a function of energy density $\epsilon$, corresponding to the parameter sets shown in the upper panel for self-bound strange and nonstrange quark matter. The gray segments of the $f/\rho_{B}$ curves represent thermodynamically unstable EOSs Fisher (1967). Lower panel: The $f/\rho_{B}$ as functions of $\mu$ and $\rho_{B}/\rho_{0}$ for strange and nonstrange quark matter.
V Stellar matter and protoquark star properties
Figure 5: The mass-radius (M-R) relations for strange and nonstrange protoquark stars corresponding to the same parameter sets shown in Fig. 4.
@@ V Stellar matter and protoquark star properties
According to the Bodmer-Witten hypothesis Bodmer (1971); Witten (1984), absolutely stable quark matter must have an energy per baryon lower than that of the $^{56}\mathrm{Fe}$ nucleus at zero pressure and zero temperature. Accordingly, for self-bound strange quark matter, the condition $(E/A)_{\mathrm{uds}}\leq 930~\mathrm{MeV}$ must be satisfied for the hypothesis of absolutely stable strange matter to hold. In contrast, bulk nonstrange quark matter should have an energy per baryon higher than that of the confined phase, i.e., $(E/A)_{\mathrm{ud}}\geq 930~\mathrm{MeV}$. For absolutely stable nonstrange quark matter, the stability condition is reversed: $(E/A)_{\mathrm{ud}}\leq 930~\mathrm{MeV}\leq(E/A)_{\mathrm{uds}}$. In our previous work using the modified NJL model with exchange interactions Yuan et al. (2022), we demonstrated that there is sufficient parameter space to support both absolutely stable nonstrange and strange quark matter at zero temperature. However, at finite temperature, these conditions are not necessarily required to be satisfied.
In the following, we study the influence of the exchange interaction parameter $\alpha$, lepton fraction $Y_{l}$, and temperature $T$ on the interacting quark matter EOS and the structure of protoquark stars. As shown in Sec. IV, increasing the strength of the exchange channel $\alpha$ alters the nature of the chiral phase transition from first-order to a crossover. Since the relationship between chiral symmetry restoration and deconfinement remains unclear, in this work, we assume that deconfinement occurs simultaneously with the chiral phase transition. Therefore, to effectively mimic confinement at low densities, a finite vacuum pressure is introduced. According to the properties of the chiral phase transition, we adopt different prescriptions for defining the vacuum pressure B-B:
  • For a first-order phase transition at low temperature and weak vector interaction, the chiral symmetry breaking phase cannot be reliably described by the NJL model. Therefore, we consider only the chiral restored quark phase after the transition and choose the minimum vacuum pressure $B_{0}$ such that the pressure vanishes at a finite energy density, as required for modeling self-bound quark stars. Any other choice of $B$ should be larger than $B_{0}$. For example, in the upper panel of Fig. 4, for the case of $\alpha=0.5$, $Y_{l}=0.1$, and $T=40~\mathrm{MeV}$, the chiral phase transition is of first order, resulting in a discontinuity in the energy density of the EOS. To ensure that the quark star is composed of chirally restored quark matter, the $B^{1/4}$ must exceed $133~\mathrm{MeV}$, allowing the quark matter EOS to reach zero pressure at a finite energy density—thus ensuring the star is self-bound at the surface.
  • For a crossover chiral transition occurring at relatively high temperatures and with strong vector interactions in the exchange channel, the transition becomes smooth. In this case, although the choice of vacuum pressure is more flexible, selecting a sufficiently small vacuum pressure ensures that the EOS yields a finite energy density at zero pressure; nevertheless, a minimum value of $B$ is still required to guarantee that the surface baryon density $\rho_{B}$ exceeds the nuclear saturation density $\rho_{0}$. The results below will indicate that the vacuum pressure significantly influences the structure of protoquark stars, suggesting that tighter constraints in the future on this parameter are essential for a deeper understanding of the quark matter EOS and compact star configurations.
V.1 Strange stellar quark matter and stars
@@
In the left panel of Fig. 4, we present the EOS at finite temperature with neutrino trapping, and the speed of sound $c_{s}^{2}$ for stellar matter, along with the behavior of the free energy per baryon $f/\rho_{B}$ for protoquark stars. The full EOSs represent results derived from the NJL model, including the chirally broken phase. To model self-bound quark matter that can form quark stars, a vacuum pressure $B$ is added to ensure the system reaches zero pressure at a finite density. For the case of $\alpha=0.5$, $Y_{l}=0.1$, and $T=10\,\mathrm{MeV}$, a discontinuous jump in energy density appears, characteristic of a first-order phase transition. In this scenario, the vacuum pressure must be at least $B_{0}^{1/4}=133\,\mathrm{MeV}$ to ensure that the pressure vanishes at the surface of the compact star. However, when the chiral phase transition becomes a crossover, the required vacuum pressure $B$ can be smaller than $B_{0}$.